OpenStudy (anonymous):

Andy has $14 more than jack. Together they have $56. How much money does each boy have?

OpenStudy (anonymous):

let jack=x andy=jack+14 but jack+andy=56 jack+jack+14=56 2*jack=56-14=42 jack=42/2=21 andy=jack+14=21+14=35